OpenStudy (anonymous):

If 1000 square centimeters of material is available to make a box with a square base and an open top, find the largest possible volume of the box.

OpenStudy (anonymous):

\[1000=2x^2+4xy\] This is your equation you use to solve for variables, you are maximizing volume which is \[V=x^2y\] and we solve for one variable in the first equation \[\frac{ (500-x^2) }{ 2x}=y\] Substitute \[(\frac{ (500-x^2) }{ 2x})x^2\] Which simplifies to \[500x-x^3\] Take the derivative of this to be able to discern min's and max's \[\[500-3x^2=0\]\] Set equal to zero x=0 at 12.9099 Use the First derivative test to prove it is a max and you have your x value. Plug the x's back in to find your y value

OpenStudy (anonymous):

Refer to the attached solution using Mathematica 8 Home Edition.